Solve each exponential equation. Give the exact value for x.
8x = 1/64
7x = 10
4x = 2.5

4x = 5
The x's should be squared





Answer by Alan3354(69443)

Solve each exponential equation. Give the exact value for x.
-----------
Use ^(Shift 6) for exponents.
eg, 8^x = 1/64 (if that's what you meant)
8^x = 1/(8^2) = 8^-2
x = -2
========================
8x = 1/64
7x = 10
4x = 2.5

4x = 5
The x's should be squared ? What does that mean?